OpenStudy (anonymous):

Justine earned $18,000 during the first year of her job at city hall. After each year she received a 3% raise. Find her total earnings during the first five years on the job.

OpenStudy (linyu):

(18000)[5+4(3/100)+3(3/100)^2+2(3/100)^3+(3/100)^4)] i think

OpenStudy (linyu):

I'm sorry i think i must have mistaken

OpenStudy (anonymous):

oops I messed up that equation

OpenStudy (anonymous):

sooo Final=18,000(0.3)^5 isss Final= 43.74? @JoyLin

OpenStudy (anonymous):

It's alright :) @Linyu

OpenStudy (anonymous):

Oh wait. Yea. Nvm. I think mine was right.

OpenStudy (anonymous):

Oh, okay @JoyLin sweet ^^ Thank you~

OpenStudy (linyu):

but 43.74 is way less than 18000

OpenStudy (anonymous):

@Linyu I just realized that. I think it would be 1.03 instead of .03? Orr?

OpenStudy (anonymous):

wait, that still doesn't make sense. dang it. haha

OpenStudy (anonymous):

You could always do it long hand to double check. 1.) 18000*.03=x x+1800=a 2.) a*.03=b b+a=c 3.) c*.03=d d+c=e 4.) e*.03=f f+e=f 5.) f*.03=g g+f=h a+c+e+f+h=amount earned

OpenStudy (anonymous):

That's how you do it long hand.

OpenStudy (anonymous):

I'm sorry, I don't understand what the final answer would be.

OpenStudy (anonymous):

What's 18000*.03?

OpenStudy (anonymous):

540

OpenStudy (anonymous):

now what's 540+18000?

OpenStudy (anonymous):

18540

OpenStudy (anonymous):

Now look back at the list I posted. Check out #1. 18540=a Do you get what I was doing?

OpenStudy (anonymous):

The variables don't represent some crazy equation. Those are left for you to fill out because I want to make sure you understand how to do it.

OpenStudy (anonymous):

OHHH. So 18540*0.03= 556.2 b= 19,096.2 and on and on

OpenStudy (anonymous):

Yuuuup! ^.^ That's all. Sorry if I confused you!

OpenStudy (anonymous):

I know. Letters make math look scary. xD

OpenStudy (anonymous):

No, it's all good! Thank you so much for the help :D haha, it does :P

OpenStudy (anonymous):

Hopefully I've helped! :) If you have any other questions that confuse you, feel free to message me or tag me. Until then, good luck and good job! Keep up the great work!

OpenStudy (linyu):

it's 18000(1+3/100)^5 to be exact

OpenStudy (anonymous):

^there's the formula way

OpenStudy (anonymous):

Both work.

OpenStudy (linyu):

trying to challenge my eyes not to close during midnight so sorry T^T

OpenStudy (anonymous):

Sooo, $98,431.37 is the final answer? I feel like thats wrong haha. It's okay! Thanks for the help :)

OpenStudy (linyu):

woah! too big!

OpenStudy (anonymous):

I got $20,866.93 for the equation you gave me, but you have to add up all 5 years right?

OpenStudy (linyu):

it's like @JoyLin said but a slightly mistake. 18000*1.03^5 yes that's correct

OpenStudy (anonymous):

Sorry. Dx Haven't done exponential growth/decay in a long time!

OpenStudy (linyu):

@JoyLin my brain rusted on this one also

OpenStudy (anonymous):

So $20,866.93 is correct?

OpenStudy (linyu):

yes

OpenStudy (anonymous):

YAY. THANK YOU.

OpenStudy (anonymous):

Sorry that took so long! Dx

OpenStudy (anonymous):

@JoyLin It's not a problem! I learned something :D

OpenStudy (anonymous):

That's good! :) Well that's what real learning is. Asking questions and making mistakes. xD